# Compound Mutations in the Abl1 Kinase Cause Inhibitor Resistance by Shifting DFG Flip Mechanisms and Relative State Populations Gabriel Monteiro da Silva

### v1

@@ Introduction
Protein kinases (PKs) are the master regulators of eukaryotic life [Manning2002]. Through their phosphotransfer activity, kinases can induce or suppress a wide range of cellular processes, from proliferation to apoptosis [Cipak2022]. Unsurprisingly, given their central role in organisms, the activity of PKs must be tightly regulated to maintain or achieve homeostasis and their dysregulation often leads to severe disease [Cipak2022, Greenman2007, Pavlovsky2019, Greuber2013, Shah2002]. The role of the Abl1 protein tyrosine kinase in cancers caused by the Philadelphia chromosome phenotype is perhaps the most widely-known example of kinase-mediated oncogenesis [DeBraekeleer2011]. Normally, Abl1 plays a crucial role in various cellular processes, such as cell division and differentiation [Wang2014]. However, oncogenesis manifests when the gene that codes for Abl1 is disturbed such as by the reciprocal translocation observed in the Philadelphia chromosome phenotype, which results from a fusion of the BCR and Abl1 genes [DeBraekeleer2011, Pavlovsky2019, Piedimonte2019]. The protein product that results from this fusion is a constitutively-active form of Abl1, disrupting homeostasis and leading to disease [Kang2016]. Importantly, the presence of the BCR-Abl1 fusion gene is a hallmark of chronic myeloid leukemia and other forms of cancer [Kang2016].
@@
In many organisms, kinase regulation is achieved through interlocking and often redundant mechanisms [Tong2014, Xie2020]. One example of such a mechanism is the steric blocking of ATP-binding sites, in which kinases such as the Abl1 tyrosine kinase can have their active sites blocked by binding partners or by their own mobile segments, rendering them unable to bind sugar phosphates and inhibiting phosphorylation [Meng2015, Shan2009, Lakkaniga2019]. Abl1 inactivation due to its own mobile segment shifting to block its own active site is a textbook example of inhibition by intrinsic dynamics [Shukla2023].
@@
Broadly, Abl1 and other kinases are capable of self-inhibition via intrinsic dynamics through different processes, perhaps the most well-studied of which is the conformational change known as the DFG flip [Meng2015, Shan2009, Lakkaniga2019]. Aptly named, the DFG flip is characterized by significant torsions in the backbone of residues 381-388 in Abl1 [Meng2015], leading to a complete inversion of the orientations of the side chain atoms of Asp381 (D) and Phe382 (F) in the plane formed with the side chain of Ala380 (Figure 1), and similar side chain inversion of downstream A-loop residues 383 to 388. When completed, this inversion inserts Phe382 into Abl1’s ATP-binding site (previously occupied by the catalytic Asp381), inhibiting the enzyme.
@@
Elucidating the structural consequences of this conformational change was crucial for furthering our understanding of how the different states occupied by kinases relate to their affinity for inhibitors, which is of continuing importance for minimizing the broad public health burden of oncogenic kinases [Wilson2015]. This burden is made worse due to the challenges in designing reliable kinase inhibitors due to the high level of structural similarity across the human kinome [Hanson2019, Shah2018, Modi2021]. Since most kinases share a highly conserved catalytic core, strong off-target interactions can be exceedingly common, with undesirable effects [Hanson2019]. This prevalence of off-target effects is what made the discovery of Imatinib, a potent protein kinase inhibitor with extremely high selectivity for Abl1 [Cohen2021], such an important breakthrough. Imatinib’s strong preference for Abl1 puzzled researchers for years, as even structural data of Imatinib-bound complexes did not explain why Imatinib’s affinity for Abl1 was significantly higher than for related kinases such as Src [Manley2002].
@@
This apparent paradox was solved when dynamics were taken into account [Wilson2015]. Specifically, researchers found through nuclear magnetic resonance (NMR) experiments that Abl1 occupied structural conformations that interacted strongly with Imatinib far more frequently than other kinases [Wilson2015]. Specifically, Imatinib binds preferentially to the DFG out (inactive) conformation of kinases, and Abl1 is known to occupy this conformation significantly more frequently than Src or other kinases [Wilson2015]. Additionally, Imatinib induces binding-favorable conformations in Abl1 more readily than in other kinases [Wilson2015, Ayaz2023].
@@
The mechanism of Imatinib’s selectivity represents a cautionary tale on the importance of accounting for conformational dynamics when studying protein systems. While researchers have routinely used computational simulations to explore the binding energies of kinase inhibitors against previously-known kinase conformations, fewer studies have sampled the entire inactivation pathway and other intrinsic conformational changes in Abl1 [Thomas2021]. Although unfortunate given the importance of studying intrinsic kinase dynamics for properly predicting the binding of inhibitors, this research gap is not surprising. Biophysical studies have found that the energy barrier associated with Abl1 inactivation exceeds 30 kcal/mol, representing a free energy barrier that is too steep to be routinely scaled within the timescales accessible with traditional simulation methods such as molecular dynamics [Xie2020]. Accordingly, previous computational studies seeking to understand the DFG flip have relied either on engineered mutations and/or accelerated dynamics methods that use biasing forces to reach the necessary timescales for the flip [Lakkaniga2019, Oruganti2021, Narayan2020, Shan2009]. Although these studies represent major contributions to the field, the application of these methods comes with substantial caveats that limit the mechanistic insights that can be drawn from the resulting dynamics, as they might lead to biasing of pathways, inaccurate reporting of transition state frequencies/energies, or improper sampling due to hysteresis [Bussi2020].
@@
Despite sampling challenges, fully understanding the details of the Abl1 inactivation pathway is of vital importance for drug discovery [Liu2006]. Atomistic simulations of Abl1 inactivation without biasing forces would provide an unparalleled wealth of information regarding the enzyme, enabling the identification of allosteric networks and potentially druggable intermediate conformations, generalization of mechanisms, and comparisons with variants or other kinases. Besides these practical motivations, sampling the Abl1 inactivation pathway would additionally be extremely valuable for basic biophysics, as it would further our understanding of enzymatic self-inhibition through intrinsic dynamics.
@@
In addition to exploring the dynamics of the wild-type enzyme, simulations could also be useful for solving longstanding controversies regarding how compound mutants grant Abl1 resistance to inhibitors [Hoemberger2020, Yamamoto2004, Gibbons2011]. Previous studies have found that variants such as Abl1 Glu255Val + Thr315Ile have significantly different activity and binding profiles than wild-type Abl1 [Zabriskie2014, Hoemberger2020]. Crucially, changes in the intrinsic dynamics of Abl1 in response to compound mutations could explain many of the complex effects observed in inhibitor-resistant Abl1 variants.
@@
Beyond the direct impacts of such a study on our understanding of Abl1 kinase and its pharmacology, insights obtained from exploring the molecular mechanisms and propensity for undergoing spontaneous inactivation are potentially transferable to other kinases carrying the DFG motif. Since the DFG motif is nearly universally conserved within the kinome [Modi2021], this represents significant opportunities for drug discovery and basic biomedicine. Further, other enzymes such as phosphatases carry conserved structural elements such as the WPD motif that have been previously observed to regulate function through backbone torsion mechanisms [Yeh2023] that are analogous to the DFG flip. Insights obtained from our simulations could also be transferable to these systems.
@@
With this motivation, we have set out to systematically simulate and analyze the dynamics of Abl1 inactivation, specifically those involving the inactivation pathway represented by the DFG flip. To achieve this, we have used the Weighted Ensemble (WE) enhanced sampling approach [Russo2022, Zwier2015, Zuckerman2017]. As the name suggests, enhanced sampling methods are strategies designed to significantly increase the timescales accessible in atomistic simulations, allowing for the effective sampling of physiologically-relevant rare events such as the binding of ligands and large conformational rearrangements [Henin2022].
@@
In this work, we report thoroughly sampled DFG flip inactivation pathways for a wild-type tyrosine kinase using simulation methods without biasing forces. Through an aggregate simulation time of over 20 s, we have collected dozens of uncorrelated and unbiased inactivation events, starting from the lowest energy conformation of the Abl1 kinase core (PDB 6XR6) [Xie2020]. Our results provided us with sufficient data to put forth a broad mechanism for the DFG flip that synthesizes observations made by many previous works regarding kinase structural elements that relate to the DFG flip.
@@
While previous studies have attributed the inhibitor resistance phenotype of Abl1 Glu255Val/Lys Thr315Ile to steric clashes with would-be inhibitors [Gibbons2011, Yamamoto2004], our results reveal that the mechanism for resistance is considerably more complex. Specifically, we have found that the aforementioned variants dramatically shift the conformational preference of apo Abl1, reducing its propensity for undergoing spontaneous inactivation through the DFG flip pathway. This phenotype translates into a major reduction of the population of the DFG-out (inactive) state, the kinase conformation to which Type-II inhibitors preferentially bind. Thus, it follows that drug resistance in the variants we studied stems not just from steric clashes with would-be inhibitors, but also from reduced inhibitor affinity during the conformational search step of binding.
@@
Our results moreover inform predictions about the propensity for the DFG flip in other kinases. By identifying the main electrostatic interactions driving spontaneous inactivation in Abl1, we were able to predict how these interactions might change in Abl1’s orthologs including Src, which is known to undergo spontaneous inactivation at much lower rates than Abl1 [Wilson2015]. Crucially, the changes in the N-Lobe negative charge density from Src to Abl1 strongly hint at the atomistic origins of these kinases’ contrasting conformational landscapes, as we found that repulsive interactions driven by negatively-charged N-Lobe residues are important for driving Abl1’s spontaneous inactivation.
@@
Considering the steep free energy barrier of roughly 32 kcal/mol associated with the DFG-in to DFG-out transition in the wild-type Abl1 kinase [Xie2020], and previous uses of the Weighted Ensemble simulation method for sampling transitions with comparable energy barriers, [Saglam2019] we elected to study the DFG flip in the wild-type Abl1 kinase core (PDB 6XR6, [Xie2020]) using the WE methodology. Applying WE provided us with an abundance of data that enabled us to derive a coherent mechanism for different pathways spanning the transition, and their accompanying probabilities. As a supervised enhanced sampling method, WE employs progress coordinates (PCs) to track the time-dependent evolution of a system from one or more basis states towards a target state. Although weighted ensemble simulations are unbiased in the sense that no biasing forces are added over the course of the simulations, the selection of progress coordinates and the bin definitions can potentially bias the results towards specific pathways [Zuckerman2017].
@@
In our selection of progress coordinates, we were mindful of this caveat and chose progress coordinates that enhance the sampling while allowing for a variety of different pathways. First, we defined PC1 as the distance between the polar hydrogen of protonated Asp381 and the backbone oxygen of Val299. This choice of progress coordinate is rooted in the observation that the formation of a strong hydrogen interaction between this pair of atoms is likely an important step in the DFG flip in Abl1 kinase [Shan2009]. For PC2, we chose the distance between the gamma carbon of Phe382 and the gamma carbon of Tyr253, which is substantially reduced in DFG-out conformations because the Phe382 side chain approaches the Tyr253 side chain after the flip. PC2 tracks the Phe382 flip, and is pathway-agnostic for that particular event, as the distance between the atom pairs should decrease independently of Phe382 flipping clockwise (passing Thr315) or counterclockwise (passing His361). Finally, we chose the angle formed by Phe382’s gamma carbon, Asp381’s protonated side chain oxygen (OD2), and Lys378’s backbone oxygen as PC3 based on observations from a study that used a similar PC to sample the DFG flip in Aurora Kinase B using metadynamics [Lakkaniga2019]. This angular PC3 should increase or decrease (based on the pathway) during the DFG flip, with peak differences at intermediate DFG configurations, and then revert to its initial state when the flip concludes. Since both an increase or decrease of PC3 accurately tracks flipping events, this progress coordinate is convenient for directing the simulations towards our DFG-out target without imposing significant pathway biases. Our progress coordinates are further illustrated in Figure 2.
@@
Having chosen our set of PCs, we ran WE simulations of wild-type Abl1 starting in the DFG-in conformation with the DFG-out conformation as a target state. Using the apo Abl1 DFG-in and DFG-out structures as references (6XR6 and 6XR7, respectively), we defined bins corresponding to the initial and target states, then used the minimal, adaptive binning (MAB) scheme to further discretize the PC space between the initial and target bins [Torrillo2021]. The MAB scheme expedites sampling by automatically distributing bins on the fly based on the evolution of the system, keeping the usage of computational resources close to constant and maximizing the sampling at bottleneck and boundary PC regions. We ran our WE simulations for 200 iterations, with 8 walkers per bin, 297 bins, totaling a maximum of 2376 segments per iteration, and an iteration time (tau) of 50 picoseconds. Each WE run was repeated two times, totalling three replicates.
@@
Using these parameters, we observed the first DFG flip event in wild-type Abl1 in the first replicate after roughly 157 iterations. Analysis of the concatenated trajectory for this crossing event revealed the existence of putative intermediate states associated with DFG torsions. For clarity, we will refer to these putative states as S1 or DFG-inter. Walkers continued to cross to the DFG-out state until the end of the simulations. To deconvolute the abundance of high-dimensional data obtained from our WE simulations, we ran a variety of analysis methods to measure the relative probability accumulated for walkers in the initial (DFG-in), putative intermediate (S1 or DFG-inter), and target (DFG-out) states. In addition to measuring walker probabilities in each state, we also used the LPATH method [Bogetti2023] to track the evolution of all successful walkers (those that made it to the DFG-out state) and recovered their pathways and associated probabilities. The results for this analysis are summarized in Figure 3.
@@
LPATH revealed two predominant pathways for the DFG flip that branch-off at the DFG-inter to DFG-out transition. Specifically, while Asp381 always flips clockwise towards Val299 in all sampled transitions, the flipping of Phe382 is more plastic as it can either flip counter-clockwise after Asp381 nearly completes its flip (staggered DFG flip, Pathway 1) or flip clockwise together with Asp381 (concerted DFG flip, Pathway 2). From our analysis, Pathway 2 has a significantly higher probability than Pathway 1, which is intuitive given that it represents a shorter path to the DFG-out state in terms of the required backbone torsions.
@@
As expected given the significant energy barrier for the flips, the relative probability of walkers in the DFG-out state is extremely low (roughly 1.9 ), while the DFG-inter state is more often occupied (with a probability of 1.3 ). These relative probabilities qualitatively agree with the large expected free energy barrier for the DFG-in to DFG-out transition (3̃2 kcal/mol), and with our observation of a putative metastable DFG-inter state that is missed by NMR experiments due to its low occupancy.
@@
Making use of the wealth of structural information generated from our weighted ensemble simulations of the DFG flip in Abl1, we sought to determine a general molecular mechanism involving N-Lobe structural elements for the DFG flip pathways sampled in our simulations. To do so, we selected sample trajectories containing the successful flip event (DFG-in to DFG-out) along the most frequent pathway (concerted flip), and analyzed the evolution of key structural elements thought or known to be important for the DFG flip. The results of this analysis are reported in Figure 4, and the selected trajectories accompanied by their topology files are available in the project’s GitHub repository [abl1_dfgflip_westpa_rawdata].
In the concerted pathway, the first stage of the flip corresponds to the partial flips of Asp381 and Phe382 (DFG-in to DFG-inter), and the second corresponds to the completion of the Asp381 flip in concert with the completion of the Phe382 flip (DFG-inter to DFG-out). In the former, the side chain of Asp381 makes its way into the vicinity of the initial position of Phe382’s side chain through a relatively slow process (spanning 5 ns in the selected trajectory). First, P-Loop fluctuations break the interaction between the Asp381 polar hydrogen and the P-Loop Tyr253’s oxygen, driving Asp381 towards the solvent-accessible gap directly below Thr315. Upon entering that gap, the polar hydrogen of protonated Asp381 forms a brief hydrogen interaction with Thr315’s side chain, which is a weak hydrogen acceptor in this context. The formation of this interaction immediately provides insights into the effects of Thr315 mutations on the propensity for the DFG flip. By analyzing the distributions of the distances between the atoms involved in the Thr315-Asp381 interaction in our WE simulations during DFG torsions, we observed a strong correlation between the strength of the Thr315-Asp381 interaction and DFG conformations (Supplementary Figure 1).
Next, the side chain of Asp381 crosses over towards Phe382’s initial position, forming a stable hydrogen bond with the backbone oxygen of Val299. This interaction is known to be important for the flip, and was previously described as a mechanism for the significantly increased propensity for the DFG flip in protonated Asp381 [Shan2009]. The formation of this interaction marks the end of the first step of the proposed mechanism for the concerted DFG flip pathway, as now the side chains of Asp381 and Phe382 shifted from their original positions and both point towards the C-Helix side of the N-Lobe (DFG-inter). Importantly, since the side chain of Asp381 in this DFG-inter conformation points away from the ATP-binding site, it is unlikely that this conformation is capable of productive phosphorylation.
@@
Previous studies simulating the DFG flip have found that brief breaks in the salt bridge formed by Lys271 and Glu286 coincide with Asp381 and Phe382 flipping [Narayan2020]. Overwhelmingly conserved among kinase domains [Modi2021], the Lys271-Glu286 interaction plays a significant role in stabilizing the N-Lobe architecture by anchoring the C-Helix. The correlation between Lys271-Glu286 salt bridge occupancy and DFG torsions was also captured in concerted DFG flips from our WE simulations, as the distances between the atoms involved in forming the salt bridge are significantly larger in intermediate DFG conformations than they are in the DFG-in or DFG-out states (Figure 5). These results provide further evidence for the role of the Lys271-Glu286 salt bridge in the DFG flip mechanism.
@@
As a crucial caveat, a clockwise flip of Phe382 (in the concerted flip pathway) puts its side chain in the vicinity of that of His361, which could potentially stabilize it in an intermediate conformation through - stacking interactions. Indeed, the angles and distances of the Phe382-His361 side chain observed in trajectories where Phe382 flips clockwise could be indicative of - stacking. The approximations inherent to the force field used in this simulation (CHARMM36M, [Huang2016]) prevent us from simulating these interactions directly, and could lead to this pathway being underrepresented in the resulting ensemble.

### v2

@@ Introduction
Protein kinases (PKs) are the master regulators of eukaryotic life [1]. Through their phosphotransfer activity, kinases can induce or suppress a wide range of cellular processes, from proliferation to apoptosis [2]. Unsurprisingly, given their central role in organisms, the activity of PKs must be tightly regulated to maintain or achieve homeostasis and their dysregulation often leads to severe disease [2, 3, 4, 5, 6]. The role of the Abl1 protein tyrosine kinase in cancers caused by the Philadelphia chromosome phenotype is perhaps the most widely-known example of kinase-mediated oncogenesis [7]. Normally, Abl1 plays a crucial role in various cellular processes, such as cell division and differentiation [8]. However, oncogenesis manifests when the gene that codes for Abl1 is disturbed such as by the reciprocal translocation observed in the Philadelphia chromosome phenotype, which results from a fusion of the BCR and Abl1 genes [7, 4, 9]. The protein product that results from this fusion is a constitutively-active form of Abl1, disrupting homeostasis and leading to disease [10]. Importantly, the presence of the BCR-Abl1 fusion gene is a hallmark of chronic myeloid leukemia and other forms of cancer [10].
@@
In many organisms, kinase regulation is achieved through interlocking and often redundant mechanisms [11, 12]. One example of such a mechanism is the steric blocking of ATP-binding sites, in which kinases such as the Abl1 tyrosine kinase can have their active sites blocked by binding partners or by their own mobile segments, rendering them unable to bind sugar phosphates and inhibiting phosphorylation [13, 14, 15]. Abl1 inactivation due to its own mobile segment shifting to block its own active site is a textbook example of inhibition by intrinsic dynamics [16].
@@
Broadly, Abl1 and other kinases are capable of self-inhibition via intrinsic dynamics through different processes, perhaps the most well-studied of which is the conformational change known as the DFG flip [13, 14, 15]. Aptly named, the DFG flip is characterized by significant torsions in the backbone of residues 381-388 in Abl1 [13], leading to a complete inversion of the orientations of the side chain atoms of Asp381 (D) and Phe382 (F) in the plane formed with the side chain of Ala380 (Figure 1), and similar side chain inversion of downstream A-loop residues 383 to 388. When completed, this inversion inserts Phe382 into Abl1’s ATP-binding site (previously occupied by the catalytic Asp381), inhibiting the enzyme.
@@
Elucidating the structural consequences of this conformational change was crucial for furthering our understanding of how the different states occupied by kinases relate to their affinity for inhibitors, which is of continuing importance for minimizing the broad public health burden of oncogenic kinases [17]. This burden is made worse due to the challenges in designing reliable kinase inhibitors due to the high level of structural similarity across the human kinome [18, 19, 20]. Since most kinases share a highly conserved catalytic core, strong off-target interactions can be exceedingly common, with undesirable effects [18]. This prevalence of off-target effects is what made the discovery of Imatinib, a potent protein kinase inhibitor with extremely high selectivity for Abl1 [21], such an important breakthrough. Imatinib’s strong preference for Abl1 puzzled researchers for years, as even structural data of Imatinib-bound complexes did not explain why Imatinib’s affinity for Abl1 was significantly higher than for related kinases such as Src [22].
@@
This apparent paradox was solved when dynamics were taken into account [17]. Specifically, researchers found through nuclear magnetic resonance (NMR) experiments that Abl1 occupied structural conformations that interacted strongly with Imatinib far more frequently than other kinases [17]. Specifically, Imatinib binds preferentially to the DFG out (inactive) conformation of kinases, and Abl1 is known to occupy this conformation significantly more frequently than Src or other kinases [17]. Additionally, Imatinib induces binding-favorable conformations in Abl1 more readily than in other kinases [17, 23].
@@
The mechanism of Imatinib’s selectivity represents a cautionary tale on the importance of accounting for conformational dynamics when studying protein systems. While researchers have routinely used computational simulations to explore the binding energies of kinase inhibitors against previously-known kinase conformations, fewer studies have sampled the entire inactivation pathway and other intrinsic conformational changes in Abl1 [24]. Although unfortunate given the importance of studying intrinsic kinase dynamics for properly predicting the binding of inhibitors, this research gap is not surprising. Biophysical studies have found that the energy barrier associated with Abl1 inactivation exceeds 30 kcal/mol, representing a free energy barrier that is too steep to be routinely scaled within the timescales accessible with traditional simulation methods such as molecular dynamics [12]. Accordingly, previous computational studies seeking to understand the DFG flip have relied either on engineered mutations and/or accelerated dynamics methods that use biasing forces to reach the necessary timescales for the flip [15, 25, 26, 14]. Although these studies represent major contributions to the field, the application of these methods comes with substantial caveats that limit the mechanistic insights that can be drawn from the resulting dynamics, as they might lead to biasing of pathways, inaccurate reporting of transition state frequencies/energies, or improper sampling due to hysteresis [27].
@@
Despite sampling challenges, fully understanding the details of the Abl1 inactivation pathway is of vital importance for drug discovery [28]. Atomistic simulations of Abl1 inactivation without biasing forces would provide an unparalleled wealth of information regarding the enzyme, enabling the identification of allosteric networks and potentially druggable intermediate conformations, generalization of mechanisms, and comparisons with variants or other kinases. Besides these practical motivations, sampling the Abl1 inactivation pathway would additionally be extremely valuable for basic biophysics, as it would further our understanding of enzymatic self-inhibition through intrinsic dynamics.
@@
In addition to exploring the dynamics of the wild-type enzyme, simulations could also be useful for solving longstanding controversies regarding how compound mutants grant Abl1 resistance to inhibitors [29, 30, 31]. Previous studies have found that variants such as Abl1 Glu255Val + Thr315Ile have significantly different activity and binding profiles than wild-type Abl1 [32, 29]. Crucially, changes in the intrinsic dynamics of Abl1 in response to compound mutations could explain many of the complex effects observed in inhibitor-resistant Abl1 variants.
@@
Beyond the direct impacts of such a study on our understanding of Abl1 kinase and its pharmacology, insights obtained from exploring the molecular mechanisms and propensity for undergoing spontaneous inactivation are potentially transferable to other kinases carrying the DFG motif. Since the DFG motif is nearly universally conserved within the kinome [20], this represents significant opportunities for drug discovery and basic biomedicine. Further, other enzymes such as phosphatases carry conserved structural elements such as the WPD motif that have been previously observed to regulate function through backbone torsion mechanisms [33] that are analogous to the DFG flip. Insights obtained from our simulations could also be transferable to these systems.
@@
With this motivation, we have set out to systematically simulate and analyze the dynamics of Abl1 inactivation, specifically those involving the inactivation pathway represented by the DFG flip. To achieve this, we have used the Weighted Ensemble (WE) enhanced sampling approach [34, 35, 36]. As the name suggests, enhanced sampling methods are strategies designed to significantly increase the timescales accessible in atomistic simulations, allowing for the effective sampling of physiologically-relevant rare events such as the binding of ligands and large conformational rearrangements [37].
@@
In this work, we report thoroughly sampled DFG flip inactivation pathways for a wild-type tyrosine kinase using simulation methods without biasing forces. Through an aggregate simulation time of over 20 s, we have collected dozens of uncorrelated and unbiased inactivation events, starting from the lowest energy conformation of the Abl1 kinase core (PDB 6XR6) [12]. Our results provided us with sufficient data to put forth a broad mechanism for the DFG flip that synthesizes observations made by many previous works regarding kinase structural elements that relate to the DFG flip.
@@
While previous studies have attributed the inhibitor-resistant phenotype of Abl1 Glu255Val/Lys Thr315Ile to steric clashes with would-be inhibitors [31, 30], our results reveal that the mechanism for resistance is considerably more complex. Specifically, we have found that the aforementioned variants dramatically shift the conformational preference of apo Abl1, reducing its propensity for undergoing spontaneous inactivation through the DFG flip pathway. This phenotype translates into a major reduction of the population of the DFG-out (inactive) state, the kinase conformation to which Type-II inhibitors preferentially bind. Thus, it follows that drug resistance in the variants we studied stems not just from steric clashes with would-be inhibitors, but also from reduced inhibitor affinity during the conformational search step of binding.
@@
Our results moreover inform predictions about the propensity for the DFG flip in other kinases. By identifying the main electrostatic interactions driving spontaneous inactivation in Abl1, we were able to predict how these interactions might change in Abl1’s orthologs including Src, which is known to undergo spontaneous inactivation at much lower rates than Abl1 [17]. Crucially, the changes in the N-Lobe negative charge density from Src to Abl1 strongly hint at the atomistic origins of these kinases’ contrasting conformational landscapes, as we found that repulsive interactions driven by negatively-charged N-Lobe residues are important for driving Abl1’s spontaneous inactivation.
@@
Considering the steep free energy barrier of roughly 32 kcal/mol associated with the DFG-in to DFG-out transition in the wild-type Abl1 kinase [12], and previous uses of the Weighted Ensemble simulation method for sampling transitions with comparable energy barriers, [38] we elected to study the DFG flip in the wild-type Abl1 kinase core (PDB 6XR6, [12]) using the WE methodology. Applying WE provided us with an abundance of data that enabled us to derive a coherent mechanism for different pathways spanning the transition, and their accompanying probabilities. As a supervised enhanced sampling method, WE employs progress coordinates (PCs) to track the time-dependent evolution of a system from one or more basis states towards a target state. Although weighted ensemble simulations are unbiased in the sense that no biasing forces are added over the course of the simulations, the selection of progress coordinates and the bin definitions can potentially bias the results towards specific pathways [36]. Additionally, traditional WEMD simulations do not explicitly enhance sampling along orthogonal degrees of freedom (those not captured by the progress coordinates). In practice, this means that insufficient PC definitions can lead to poor sampling.
@@
In our selection of progress coordinates, we were mindful of this caveat and chose progress coordinates that enhance the sampling while allowing for a variety of different pathways. First, we defined PC1 as the distance between the polar hydrogen of protonated Asp381 and the backbone oxygen of Val299. This choice of progress coordinate is rooted in the observation that the formation of a strong hydrogen interaction between this pair of atoms is likely an important step in the DFG flip in Abl1 kinase [14]. For PC2, we chose the distance between the gamma carbon of Phe382 and the gamma carbon of Tyr253, which is substantially reduced in DFG-out conformations because the Phe382 side chain approaches the Tyr253 side chain after the flip. PC2 tracks the Phe382 flip, and is pathway-agnostic for that particular event, as the distance between the atom pairs should decrease independently of Phe382 flipping clockwise (passing Thr315) or counterclockwise (passing His361). Finally, we chose the angle formed by Phe382’s gamma carbon, Asp381’s protonated side chain oxygen (OD2), and Lys378’s backbone oxygen as PC3 based on observations from a study that used a similar PC to sample the DFG flip in Aurora Kinase B using metadynamics [15]. This angular PC3 should increase or decrease (based on the pathway) during the DFG flip, with peak differences at intermediate DFG configurations, and then revert to its initial state when the flip concludes. Since both an increase or decrease of PC3 accurately tracks flipping events, this progress coordinate is convenient for directing the simulations towards our DFG-out target without imposing significant pathway biases. Our progress coordinates are further illustrated in Figure 2.
For downstream analysis, we used two pseudodihedrals previously defined in the existing Abl1 DFG flip simulation literature [13] to identify and discriminate between DFG states. The first (dihedral 1) tracks the flip state of Asp381, and is formed by the beta carbon of Ala380, the alpha carbon of Ala380, the alpha carbon of Asp381, and the gamma carbon of Asp381. The second (dihedral 2) tracks the flip state of Phe382, and is formed by the beta carbon of Ala380, the alpha carbon of Ala380, the alpha carbon of Phe381, and the gamma carbon of Phe381. These pseudodihedrals, when plotted in relation to each other, clearly distinguish between the initial DFG-in state, the target DFG-out state, and potential intermediate states in which either Asp381 or Phe381 has flipped.
@@
Having chosen our set of PCs, we ran WE simulations of wild-type Abl1 starting in the DFG-in conformation with the DFG-out conformation as a target state. Using the apo Abl1 DFG-in and DFG-out structures as references (6XR6 and 6XR7, respectively), we defined bins corresponding to the initial and target states, then used the minimal, adaptive binning (MAB) scheme to further discretize the PC space between the initial and target bins [39]. The MAB scheme expedites sampling by automatically distributing bins on the fly based on the evolution of the system, keeping the usage of computational resources close to constant and maximizing the sampling at bottleneck and boundary PC regions. We ran our WE simulations for 200 iterations, with 8 walkers per bin, 297 bins, totaling a maximum of 2376 segments per iteration, and an iteration time (tau) of 50 picoseconds. Each WE run was repeated two times, totalling three replicates.
@@
Using these parameters, we observed the first DFG flip event in wild-type Abl1 in the first replicate after roughly 157 iterations. Analysis of the concatenated trajectory for this crossing event revealed the existence of a putative intermediate state associated with DFG torsions. For clarity, we will refer to this putative state as DFG-inter. Walkers continued to cross to the DFG-out state until the end of the simulations. To deconvolute the abundance of high-dimensional data obtained from our WE simulations, we ran a variety of analysis methods to measure the relative probability accumulated for walkers in the initial (DFG-in), putative intermediate (DFG-inter), and target (DFG-out) states. The results for this analysis are summarized in Figure 3.
@@
Our analysis revealed two predominant pathways for the DFG flip that branch-off at the DFG-inter to DFG-out transition. Specifically, while Asp381 always flips clockwise towards Val299 in all sampled transitions, the flipping of Phe382 is more plastic as it can either flip counter-clockwise after Asp381 nearly completes its flip (staggered DFG flip, Pathway 1) or flip clockwise together with Asp381 (concerted DFG flip, Pathway 2). From our analysis, Pathway 2 has a significantly higher probability than Pathway 1, which is intuitive given that it represents a shorter path to the DFG-out state in terms of the required backbone torsions.
@@
As expected given the significant energy barrier for the flips, the relative probability of walkers reaching the DFG-out state is extremely low (roughly 1.9 ). These relative probabilities qualitatively agree with the large expected free energy barrier for the DFG-in to DFG-out transition (3̃2 kcal/mol).
@@
Making use of the wealth of structural information generated from our weighted ensemble simulations of the DFG flip in Abl1, we sought to determine a general molecular mechanism involving N-Lobe structural elements for the DFG flip pathways sampled in our simulations. To do so, we selected sample trajectories containing the successful flip event (DFG-in to DFG-out) along the most frequent pathway (concerted flip), and analyzed the evolution of key structural elements thought or known to be important for the DFG flip. The results of this analysis are reported in Figure 4, and the selected trajectories accompanied by their topology files are available in the project’s GitHub repository [40].
In the concerted pathway, the first stage of the flip corresponds to the partial flips of Asp381 and Phe382 (DFG-in to DFG-inter), and the second corresponds to the completion of the Asp381 flip in concert with the completion of the Phe382 flip (DFG-inter to DFG-out). In the former, the side chain of Asp381 makes its way into the vicinity of the initial position of Phe382’s side chain through a relatively slow process (spanning 5 ns in the selected trajectory). First, P-Loop fluctuations break the interaction between the Asp381 polar hydrogen and the P-Loop Tyr253’s oxygen, driving Asp381 towards the solvent-accessible gap directly below Thr315 [10.145/118455]. Upon entering that gap, the polar hydrogen of protonated Asp381 forms a brief hydrogen interaction with Thr315’s side chain, which is a weak hydrogen acceptor in this context. The formation of this interaction immediately provides insights into the effects of Thr315 mutations on the propensity for the DFG flip. By analyzing the distributions of the distances between the atoms involved in the Thr315-Asp381 interaction in our WE simulations during DFG torsions, we observed a strong correlation between the strength of the Thr315-Asp381 interaction and DFG conformations (Supplementary Figure 1).
Next, the side chain of Asp381 crosses over towards Phe382’s initial position, forming a stable hydrogen bond with the backbone oxygen of Val299. This interaction is known to be important for the flip, and was previously described as a mechanism for the significantly increased propensity for the DFG flip in protonated Asp381 [14]. The formation of this interaction marks the end of the first step of the proposed mechanism for the concerted DFG flip pathway, as now the side chains of Asp381 and Phe382 shifted from their original positions and both point towards the C-Helix side of the N-Lobe (DFG-inter). Importantly, since the side chain of Asp381 in this DFG-inter conformation points away from the ATP-binding site, it is unlikely that this conformation is capable of productive phosphorylation.
@@
Previous studies simulating the DFG flip have found that brief breaks in the salt bridge formed by Lys271 and Glu286 coincide with Asp381 and Phe382 flipping [26]. Overwhelmingly conserved among kinase domains [20], the Lys271-Glu286 interaction plays a significant role in stabilizing the N-Lobe architecture by anchoring the C-Helix. The correlation between Lys271-Glu286 salt bridge occupancy and DFG torsions was also captured in concerted DFG flips from our WE simulations, as the distances between the atoms involved in forming the salt bridge are significantly larger in intermediate DFG conformations than they are in the DFG-in or DFG-out states (Figure 5). These results provide further evidence for the role of the Lys271-Glu286 salt bridge in the DFG flip mechanism.
@@
As a crucial caveat, a clockwise flip of Phe382 (in the concerted flip pathway) puts its side chain in the vicinity of that of His361, which could potentially stabilize it in an intermediate conformation through - stacking interactions. Indeed, the angles and distances of the Phe382-His361 side chain observed in trajectories where Phe382 flips clockwise could be indicative of - stacking. The approximations inherent to the force field used in this simulation (CHARMM36M, [41]) prevent us from simulating these interactions directly, and could lead to this pathway being underrepresented in the resulting ensemble.